OpenStudy (anonymous):

Show all work. Mary purchased a package of 9 different plants, but she only needed 6 plants for planting. In how many ways can she select the 6 plants from the package to be planted?

OpenStudy (kropot72):

The number of ways of selecting r objects from a total of n, all different, is given by \[\frac{n!}{r!(n-r)!}\] The answer is found from the following: \[9C6=\frac{9!}{6!(9-6)!}=\frac{9\times 8\times 7}{3\times 2}=?\]

OpenStudy (kropot72):

@stef15 Do you understand?

OpenStudy (anonymous):

do i just times the top and the botom and then reduce and that wll be the answer

OpenStudy (anonymous):

504/6=84 is thi right

OpenStudy (kropot72):

Yes! The correct answer is that the 6 plants can be selected in 84 ways.